# Industry Expectations and Skill Demands in Quantum Software Testing

### v1

@@ 5.4. Threats to Validity
The credibility of the findings may be influenced by the reliance on job postings as data sources. Job advertisements do not always represent the full scope of responsibilities or competencies required in practice, and their descriptions may reflect aspirational or simplified portrayals. To mitigate this limitation, we included postings from multiple organizations and cross-validated responsibilities against the literature. Nevertheless, the interpretation of job texts necessarily involves researcher judgment. The transferability of our results is constrained by the sample’s geographic concentration in North America and Europe. While these regions dominate the current quantum industry, the findings may not fully capture practices emerging in other regions where quantum research is growing. Furthermore, as job postings evolve rapidly, our dataset provides a snapshot of an industry undergoing rapid transformation rather than a static representation of its requirements. Reflexivity was considered throughout the analysis. The authors’ backgrounds in software engineering and research may have influenced how testing responsibilities were interpreted. Finally, rigor was established through systematic coding and comparison across postings, triangulated with current literature on quantum software testing. Although this study is limited to publicly available data, it provides a transparent and replicable basis for understanding the evolving expectations of testing in quantum software engineering.
@@ 6. Conclusions and Future Work
This study analyzed 110 job postings to characterize how testing is defined and what skills are required in quantum software engineering. The results indicate that testing in this domain combines classical quality assurance practices with experimental validation, with emphasis on calibration, control, and hybrid quantum–classical verification. Employers seek professionals who combine programming and automation capabilities with quantum-specific knowledge, alongside collaboration, communication, and analytical reasoning across disciplinary boundaries.
The findings indicate that required skills are not static and are likely to change as quantum technologies evolve. Current roles emphasize close interaction with hardware, experimental validation, and low-level system understanding, reflecting the early stage of industrial adoption. As quantum platforms mature, testing practices are likely to shift toward higher levels of abstraction, greater automation, and more standardized verification workflows. This evolution has direct implications for curriculum design and certification frameworks, which should be designed to accommodate changing skill profiles over time. Educational programs may need to balance enduring software engineering fundamentals with emerging quantum-specific competencies, while certification frameworks should remain adaptable to technological change rather than codifying a fixed set of skills.

### v2

@@ 5.4. Threats to Validity
The credibility of the findings is influenced by reliance on job postings as data sources, since advertisements may provide incomplete, simplified, or aspirational descriptions and require researcher interpretation. This limitation was partially mitigated by sampling postings from multiple organizations and cross validating reported responsibilities with existing literature. However, some postings include activities more closely related to quantum hardware engineering, which may bias the identified testing skill set toward hardware optimization rather than software oriented verification and validation. The transferability of the results is constrained by the geographic concentration of the sample in North America and Europe and by the rapidly evolving nature of job advertisements, which capture a time specific snapshot of an emerging industry. Reflexivity was explicitly considered, as the authors’ backgrounds in software engineering and research may have influenced how testing related responsibilities were interpreted. Finally, rigor was supported through systematic coding, comparison across postings, and triangulation with current literature, providing a transparent and replicable analysis based on publicly available data.
@@ 6. Conclusions and Future Work
This study analyzed 110 job postings to characterize how testing is defined and what skills are required in quantum software engineering. The results indicate that testing in this domain combines classical quality assurance practices with experimental validation, with emphasis on calibration, control, and hybrid quantum–classical verification. Employers seek professionals who combine programming and automation capabilities with quantum-specific knowledge, alongside collaboration, communication, and analytical reasoning across disciplinary boundaries. The findings indicate that required skills are not static and are likely to change as quantum technologies evolve. Current roles emphasize close interaction with hardware, experimental validation, and low-level system understanding, reflecting the early stage of industrial adoption. As quantum platforms mature, testing practices are likely to shift toward higher levels of abstraction, greater automation, and more standardized verification workflows. This evolution has direct implications for curriculum design and certification frameworks, which should be designed to accommodate changing skill profiles over time. Educational programs may need to balance enduring software engineering fundamentals with emerging quantum-specific competencies, while certification frameworks should remain adaptable to technological change rather than codifying a fixed set of skills. This study opens several avenues for follow-up research. First, a longitudinal replication (e.g., re-collecting postings in 12–24 months using the same protocol) would allow us to quantify how required testing skills evolve as platforms mature, highlighting emerging and declining competencies and the pace at which testing moves up the abstraction stack. Second, given that the dataset already enables richer exploration, future work can conduct deeper comparative analyses across regions, company types, and role seniority to identify statistically supported differences. Third, a more formal comparison with classical software testing roles—using established competency models from software testing/QA and mapping overlaps and deltas—would clarify which aspects of quantum testing reflect continuity with traditional practice and which are genuinely distinctive. Finally, future studies should explicitly classify skills by quantum software stack and framework, distinguishing generalizable testing competencies from those that are technology- or vendor-specific; this would better inform curriculum design and certification frameworks by separating enduring fundamentals from rapidly shifting, platform-bound skills.
7. Acknowledgments
This work has been partially supported by the following projects: ”QUASAR: QUAntum software engineering for Secure, Affordable, and Reliable systems”, grant 2022T2E39C, under the PRIN 2022 MUR program funded by the EU - NGEU; SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union – NextGenerationEU; Patto territoriale ”Sistema universitario pugliese” – CUP F61B23000370006